A 500.0 mL solution of NaCN was prepared from a 12.0g sample. What is the pH of the solution? (Ka for HCN is 4.0x10^-10)
Answer #1

Date 11 Page x² + 2.5xlos - 1.225 x10-5 = on. earing a = 3.48410-3 m [OH-] = 3.48X103 Poha- log [or] 2 = -lop [3.48 810 3. PO
Date Page mularity of NaCN a mole Volume (2) mol y No (N= man = Molar mass 128 -0.244 9 mo. 49 8l mo mol = 0.245 mol [1L=1000